Syllabus

1 Administrative details

Location and time Kresge 327, MWF 2–3:10

Instructor and TA

Kevin Karplus 315B Appl. Sci. Thurs 2–3 and by appt. 459-4250 karplus@cse.ucsc.edu
Karen Glocer Kresge cafe  WF 3:15-4:15  phone? kag@cse.ucsc.edu

Tutors

You will be expected to see a tutor every other week—we have been budgeted enough tutor time to average fifteen minutes a week per student, so half an hour every two weeks is a fair share. Some of the students with a weaker background in writing may need to work with the tutors more frequently. Past experience has been that the best writers were also the most diligent about seeing the tutors (or that the students who used the tutors most ended up writing best—cause and effect are a little hard to separate here).

2 Homework Assignments

The homework assignments will be posted to the web in three formats: Postscript, dvi, and HTML.

To read the files on-line, use Netscape (or one of the other World-Wide Web browsers), and open the Uniform Resource Locator (URL) http://www.cse.ucsc.edu/~karplus/. This is the home page for Kevin Karplus, and contains pointers to several other documents, including the assignments for this course. If you have not used Netscape before, almost any of the CATS consultants in any of the computer labs can help you, because browsing the World-Wide Web has become a major form of recreation for the consultants.

The easiest to read on-line is the HTML, but the Latex2HTML translator occasionally makes mistakes, so there may be some strange behavior. Although it is possible to print directly from many World-Wide Web browsers, I’ll also provide PostScript files generated directly from the L^{A}TEX originals. These will generally produce a better looking printed copy that will take up less paper. There will be hypertext links (pointers) from the WWW pages to the PostScript files. The .dvi format is now readable on Macintoshes and UNIX machines with free software, and is much smaller to download than the Postscript.

Please read each of the assignments carefully—don’t rely on a vague memory of in-class discussion. Common problems that have come up with the assignments in the past are discussed in the assignment writeups, but still about half of the problems I see in turned in work are things that I specifically warn about.

3 Evaluation

CMPE 185 has no mid-term or final exams. You will be judged about 75% on the papers you write, 15% on the oral presentation near the end of the quarter, and 10% on in-class work, participation in discussions (both in class and electronic) peer editing, and so forth. The final paper is about a quarter of the total weight for the papers.

Everything must be turned in by the last day of class, Friday Dec 1, as I have a conference to attend all of finals week, and need to do the grading mostly on Dec 2.

4 Special guest lecturers

I may arrange to have some guest lectures this quarter. So far, I have arranged for Tracy Larrabee to give lectures on Oct 9 and 11, when I have to go to Michigan. In mid-October we will have a presentation by a reference librarian. As a courtesy to our guest lecturers, please make an extra effort to be on time for their lectures.

5 Schedule (due dates)

The following schedule lists the assignments for this quarter with due dates. If the due dates are changed, I will inform you in class. Note that many assignments overlap; this is the normal case for people who write for a living, as engineers do (30–50% of the work day).

You are required to turn in your early drafts, as well as the final copy—the early drafts allow us to see how well you edit and how well you respond to suggestions from your peers.

Save your assignments after I return them to you! I will ask you to turn them all in at the end of the quarter, so that I can write a fair evaluation of your work. Get a large envelope to keep all returned assignments in, so that you can turn in the entire portfolio at the end of the quarter.

original  revised or redo
assignment draft(s) final
0. intake survey (not graded)  Sep 22
1. job application and résumé Sep 22 Sep 25
2. in-class hiring memo  Sep 25
3. in-program documentation Oct 2 Oct 6 Oct 27
4. naive user documentation (HTML manual) Oct 13  Oct 16
5. library puzzle  Oct 20
6. poster presentations  Oct 23,25,27
7. term project proposal memo  Oct 30
8. describing an algorithm Nov 3 Nov 6
9. term project document specs  Nov 13
10. term project progress report  Nov 20
11. term project Nov 20 Dec 1
12. oral presentations  Nov 20, 22, 27, 29 Dec 1

Note: no final exam is scheduled, and the time slot (12-3 Dec 8) will not be used. All written work must be completed and turned in by Dec 1, the last day of class.

6 Reading assignments

I have collected all the reading assignments scattered through this reader into the following schedule, and have included chapters to read that are not mentioned explicitly in the individual assignments. Several of the chapters are listed repeatedly—this does not mean you need to read them repeatedly, but that you should read them the first time they appear, and review the relevant parts for the later assignments. Note that this reading is in addition to reading the material in this workbook, which should be read in order, with each section read a week or two before the corresponding assignment is due.

It is probably a good idea to read ahead, since you wil have more time at the beginning of the quarter than the end of the quarter.

Please turn in a page of notes on the required reading on each of the due dates! The notes will not be evaluated for content, but are our way of ensuring that you read the material actively, getting the most out of these somewhat expensive books.

assignment  read before
[HO91, Chapters 1, 3, 10, 11, and 13]  Sep 22
[HO91, Chapters 5, 7, 21, and 27]  Sep 29
[SJW79]  Oct 6
[HO91, Chapters 8, 9, 17, 22, 23, 29, and 30]  Oct 13
[HO91, Chapters 12, 14, 24, 25, and 26]  Oct 20
[HO91, Chapters 2, 4, 5, 12, 16, 18, 21, and 28]  Oct 27
[HO91, Chapters 15 and 19, Appendix A4]  Nov 3

I have not scheduled the reading of Chapters 6, 20, and 31–38, but you should make time to read them. Non-native speakers are particularly encouraged to read Chapters 29–38 early in the quarter.

References

[HO91] Thomas N. Huckin and Leslie A. Olsen. Technical Writing and Professional Communication for Nonnative Speakers of English. McGraw-Hill Book Company, New York, 1991.

[SJW79] William Strunk Jr. and E. B. White. The Elements of Style. Macmillan, New York, third edition, 1979.